# CLIP-Guided Backdoor Defense through Entropy-Based Poisoned Dataset Separation

### v1

@@ Appendix G More Results on Defenses Based on Clean Data
Table 14 presents the results of clean-data based defenses with 5% clean training data available. In this setting, the total number of successful defenses (defined as defenses achieving an Attack Success Rate (ASR) of 20% or lower, highlighted in green) is 23.
Table 14 shows the results for clean-data based defenses when a clean subset is split from the whole (100%) poisoned training dataset. With access to the full poisoned dataset, the defense performance improves significantly, with the number of successful defenses increasing to 35, marking a 50% improvement over the 5% clean data setting. This indicates that having access to the full poisoned dataset allows the defenses to operate more effectively, even in the absence of originally clean data.
Table 15 provides results for clean-data based defenses with a 10% poisoned training dataset. Even in this setting, where only a small portion of the dataset is clean, the total number of successful defenses increases from 23 (in the 5% clean data setting) to 25. This demonstrates that a 10% poisoned dataset can yield comparable performance to that obtained with 5% clean data, suggesting that even minimal clean data availability can significantly bolster defense effectiveness.

### v2

@@ Appendix G More Results on Defenses Based on Clean Data
Table F presents the results of clean-data based defenses with 5% clean training data available. In this setting, the total number of successful defenses (defined as defenses achieving an Attack Success Rate (ASR) of 20% or lower, highlighted in green) is 23.
Table F shows the results for clean-data based defenses when a clean subset is split from the whole (100%) poisoned training dataset. With access to the full poisoned dataset, the defense performance improves significantly, with the number of successful defenses increasing to 35, marking a 50% improvement over the 5% clean data setting. This indicates that having access to the full poisoned dataset allows the defenses to operate more effectively, even in the absence of originally clean data.
Table F provides results for clean-data based defenses with a 10% poisoned training dataset. Even in this setting, where only a small portion of the dataset is clean, the total number of successful defenses increases from 23 (in the 5% clean data setting) to 25. This demonstrates that a 10% poisoned dataset can yield comparable performance to that obtained with 5% clean data, suggesting that even minimal clean data availability can significantly bolster defense effectiveness.